Divide 59 into two parts so that one part is 7 less than five times the other part
Answer by stanbon(75887)

Divide 59 into two parts so that one part is 7 less than five times the other part
---
Equations:
x + y = 59
x = 5y-7
------
Substitute for "x" and solve for "y":
5y-7 + y = 59
6y = 66
y = 11
---
Solve for "x":
x = 5*11-7
x = 48
==============
Cheers,
Stan H.
==============